OpenStudy (anonymous):

Solve for x: = 3x x = -7 x = -13 x = 8 x = 2 @sodukogirl13

OpenStudy (anonymous):

rest of equation? what is equal to 3x?

OpenStudy (austinl):

We are missing a rather important part of the question...

OpenStudy (anonymous):

(x+16) ------ 3

OpenStudy (austinl):

\(\displaystyle \frac{x+16}{3}=3x\) We would first want to multiply both sides by 3 \(\displaystyle x+16=9x\) then get all x's on one side \(16=8x\) divide by 8 \(x=~?\)

OpenStudy (anonymous):

x=2 :)

OpenStudy (anonymous):

Im sooooo smart XD

OpenStudy (anonymous):

i hope so

OpenStudy (anonymous):

\[\frac{ x+16 }{ 3 }=3x\] \[3*\frac{ x+16 }{ 3 }=3x*3\] x+16=9x -x -x 16=8x 16/8=8x/8 2=x